OpenStudy (anonymous):

i still dont understand this :,( can someone help me to prove this identity? I give medals!

OpenStudy (anonymous):

what identity?

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

\[\cos (\frac{ \pi }{ 4 } +x) + \cos (\frac{ \pi }{ 4 } -x) = \sqrt{2} cosx\]

OpenStudy (perl):

expand the left side using cosine addiction/difference formula cos ( u + v) = cos u cos v - sin u sin v cos ( u - v) = cos u cos v + sin u sin v

OpenStudy (perl):

cos ( Pi/4 + x ) = cos (pi/4) cos x - sin(pi/4) sin x cos ( Pi/4 - x ) = cos (pi/4) cos x + sin(pi/4) sin x now add those two equations

OpenStudy (anonymous):

okay thanks I'm looking at the problems and comparing what I have already

OpenStudy (anonymous):

so what would I end up doing next? this is the part that I got stuck at

OpenStudy (anonymous):

Okay, so sorry my brain is fried looking at this all day. It looks like you're saying to use a sum and difference formula for cosine

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (anonymous):

@kittenlover731

OpenStudy (anonymous):

@zepdrix

OpenStudy (shamim):

Cos(π/4+x)=cosπ/4.cosx-sinπ/4.sinx

OpenStudy (anonymous):

huh?

OpenStudy (shamim):

Use my formula to prove ur identity

OpenStudy (anonymous):

your formula didnt show up in the post, only question marks

OpenStudy (sleepyjess):

refresh the page

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i think thats what @perl said above, she said to add those two equations

OpenStudy (anonymous):

thats where i got stuck, i dont know what to do after that

OpenStudy (anonymous):

@shamim what did you get after you added them?

OpenStudy (anonymous):

@sophiaedge

OpenStudy (shamim):

2cos pi/4.cosx

OpenStudy (anonymous):

im lost :( can you explain how you did that?

OpenStudy (anonymous):

sorry grace, we most definitely haven't gone this far in my class. we're just wrapping up on the 6 functions

OpenStudy (anonymous):

No worries, thanks for trying